The countries, X and Y, has consumer price indexes in 2015 equal to 160 (in country X) and to 120 (in country Y) The base year is 2000.

a) Which of the countries has experienced a higher inflation rate between the years 2000 and 2015.

b) Suppose the CPI of Country Y in 2016 (with base year 2000) is equal to 130 and the CPI in 2016 of country X is 170. What can we say about the rate of inflation (between the base year and 2016) and the price level in 2016 in the two countries? Which country has experienced a higher inflation rate and which country has a higher price level? Explain your answer and If you do not have enough information, what would you need?
Answer #1

a) Country X has experienced a higher rate of inflation between the years 2015 and 2000 as compared to countyc Y, Country X has the higher level of consumer price index.

b) the rate of inflation in country X assuming the price index in the base year to be 100 is = 170-100)/100 = 70% and thus the prices are 70% higher in 2016 than in 2000.

Similarly, the rate of unflainflin country Y = (130-100)/100 = 30% and the prices are thus 30% higher in 2016 than in 2000.

Thus, country X has experienced the higher rate of inflation and has the higher price level.

Thanks!